OpenStudy (anonymous):

PLEASE HELP!!!! A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is _____ feet per second.

OpenStudy (studygurl14):

Plug in 4 for t and solve for f(t). Then plug in 6 for t and solve again. Then find the average of those two.

OpenStudy (anonymous):

f(6) = -16(6)^2 -32(6) +90 -678 f(4)= -16(4)^2 -32(4) +90 -294

OpenStudy (studygurl14):

Wait, sorry. Ignore the last instruction. To find the rate of change, you'll need to find teh slope of the two points.

OpenStudy (studygurl14):

You messed up the sign of 32. It's + 32, not - 32

OpenStudy (studygurl14):

@CountryGurl15 do you understand?

OpenStudy (anonymous):

oh sorry, let me fix that =) hold on

OpenStudy (anonymous):

f(6)=-16(6)^2 +32(6) +90 -294 f(4)= -16(4)^2 +32(4) +90 -38

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

The second one is correct, but -294 is not.

OpenStudy (anonymous):

hmmm idk how i keep getting that one wrong!

OpenStudy (anonymous):

is it 186?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

Correct.

OpenStudy (studygurl14):

Now you have these points: (4, -38) and (6, 186), right? To find the rate of change, find teh slope of these two points.

OpenStudy (anonymous):

okay awesome thx so much! =) can i double check my slope with you once i get it please? =)

OpenStudy (anonymous):

Slope is 112 right, @StudyGurl14 =)

OpenStudy (studygurl14):

Correct. nice job. :)

OpenStudy (anonymous):

awesome so that's my answer right?

OpenStudy (studygurl14):

Yep.

OpenStudy (anonymous):

awesome! thx so much =)